A body covers 24m and 64m in first 4sec and in next 4 sec respectively. Find the initial speed

Grade:9

2 Answers

akshay Jaitpal
13 years ago

apply the formula So=ut+1/2at2

Aniket Patra
13 years ago

Let the initial speed be u..After covering 24m it will attain a velocity say v..Now for 24m apply s=ut+0.5ft2 and get the equation..Now again apply v=u+ft..and 64=vt+0.5ft2..And get the equations and solve to get the answer..(f is the acelaration)